OpenStudy (steve816):

Whoa, I didn't know this. So I know that 2 x 2 = 2 + 2, and I thought that the number 2 was the only number that had such property where multiplying itself is the same as adding by itself, but there are other numbers with this property!

OpenStudy (steve816):

\[\large \sqrt{3}\times \sqrt{3}\times \sqrt{3}=\sqrt{3}+\sqrt{3}+\sqrt{3}\]

OpenStudy (mathmale):

That equation is false, Steve. Multiplication and addition are VERY different operations.

OpenStudy (steve816):

@mathmale Check again, it is true.

OpenStudy (freckles):

\[\sqrt{3} \cdot \sqrt{3}=3 \\ \text{ so } \sqrt{3} \cdot \sqrt{3} \cdot \sqrt{3}=3 \sqrt{3} \\ \text{ and } \sqrt{3}+\sqrt{3}+\sqrt{3}=3 \sqrt{3}\]

OpenStudy (mathmale):

The burden of proof is on your shoulders. \[\sqrt{3}\sqrt{3}=3;~but ~\sqrt{3}+\sqrt{3}=2\sqrt{3}\]

OpenStudy (freckles):

I think he is just noticing something neat

OpenStudy (mathmale):

Better to keep an open mind, Steve, instead of refuting what a retired professor tells you about a math problem.

OpenStudy (steve816):

@mathmale That wasn't what I was saying.

OpenStudy (freckles):

and yes being that it is neat is an opinion :p

OpenStudy (mathmale):

Steve, if you are convinced that you are right, then defend your statement.

OpenStudy (steve816):

@freckles just proved my statement and proved you wrong!

OpenStudy (mathmale):

OK, Steve, have it your way. I'm here to help students who want to learn, not those who want primarily to argue. Take care, over and out.

OpenStudy (steve816):

I don't know what you are not getting mathmale, I am trying to say that root 3 multiplied by itself three times is the same as adding root 3 three times.

OpenStudy (steve816):

I never wanted to start an argument in the first place mathmale.

OpenStudy (steve816):

AND, I do want to learn.

OpenStudy (mathmale):

Perhaps Freckles will have more patience with you than I do at the moment.

OpenStudy (freckles):

we should look at finding more values a with the fact that \[(a)^n=\sum_{i=1}^{n} a\]

OpenStudy (freckles):

where n is any integer that we choose

OpenStudy (freckles):

guess i could have just wrote a^n=n*a

OpenStudy (mathmale):

freckles: can you prove that your statement is true for all n? 2*2 = 4; 2^2=4; YES 3*3 = 9; 3^3 = 27; NO

OpenStudy (loser66):

To me, it is NOT a property!! because it is true at P(2) , Like 2*2 = 2+2 from P(3) 2*2*2 \(\neq 2+2+2\) It has nothing to be called "property"

OpenStudy (freckles):

I didn't say a^n=n*a is true for all n

OpenStudy (freckles):

I said let's try to find values of a such that a^n=n*a is true for some value of n we choose

OpenStudy (mathmale):

I agree with Loser66. He just won (in spite of his name).

OpenStudy (mathmale):

Freckles: I'd go along with that latest statement of yours. How would you go about finding n for this situation?

OpenStudy (freckles):

for example steve stated another one which was 2*2=2+2

OpenStudy (freckles):

That was my original statement if you go back up :p

OpenStudy (mathmale):

I agree with you and Steven in this particular example. One example doesn't make a rule, does it?

OpenStudy (freckles):

no one claimed a^n=a*n was a rule

OpenStudy (freckles):

for all a and n

OpenStudy (mathmale):

OK. How would you wrap up this discussion? What have we discovered?

OpenStudy (freckles):

To me the whole point of this discussion in which you are trying to derail is trying to find values of a such that a^n=n*a is true for some value of n we choose

OpenStudy (freckles):

right now that is what I'm trying to do is find a solution to that equation

OpenStudy (freckles):

besides the ones given which were a=2 , n=2 a=sqrt(3) , n=3

OpenStudy (mathmale):

You may need to choose a particular value for "a," such as "a=2" and then work at solving the resulting equation for n.

OpenStudy (steve816):

So using frecle's equation, I found another solution when n = 4 \[\large (2^\frac{ 3 }{ 4 })^4=2^\frac{ 3 }{ 4 }*4\]

OpenStudy (freckles):

that one doesn't work that one side is 8 while the other isn't

OpenStudy (mathmale):

Steve: can you prove that using a calculator?

OpenStudy (steve816):

Waiit, hold on I wrote it wrong. The exponent is supposed to be 2/3

OpenStudy (mathmale):

Freckles essentially did that (prove the "equation" false by doing actual calculations).

OpenStudy (steve816):

\[\large (2^\frac{ 2 }{ 3 })^4=2^\frac{ 2 }{ 3 }*4\]

OpenStudy (freckles):

oh nice job steve

OpenStudy (freckles):

and I think I found one for n=5

OpenStudy (freckles):

oops n=6 \[a=6^\frac{1}{5} \text{ when } n=6\]

OpenStudy (freckles):

yep that seems an easy way just choosing values of n which results in a polynomial to solve in terms of a I guess we could also use non-real values

OpenStudy (steve816):

ohhh, so does this work for any n value?

OpenStudy (mathmale):

Steve: That's a good question, and one that I hope you can answer yourself. My initial response to y our post here was NO! I didn't stop to check whether there might be exceptions to that. 2*2 = 2^2 is true and is thus an exception.

OpenStudy (mathmale):

Make up your own examples and then look for a pattern.

OpenStudy (freckles):

yep \[a^n=an \\ a^n-an=0 \\ a(a^{n-1}-n)=0 \\ a=0 \text{ or } a^{n-1}-n=0 \\ a=0 \text{ or } a^{n-1}=n \\ a=0 \text{ or } a=n^{\frac{1}{n-1}}\]

OpenStudy (mhchen):

This is soooo beautiful.

zepdrix (zepdrix):

Hmm this is really cool post. So is this what we figured out?\[\large\rm \sum_{i=1}^n n^{\frac{1}{n-1}}\quad=\quad\prod_{i=1}^n n^{\frac{1}{n-1}}\]Does that look correct maybe?

OpenStudy (steve816):

One thing: how did you go from \[a^{n-1}=n\]to\[a=n ^{\frac{ 1 }{ n-1 }}\]

zepdrix (zepdrix):

To undo (n-1) exponent, we take the (n-1)th root of each side, ya?

OpenStudy (steve816):

ohhh, ok I see!

OpenStudy (kainui):

Is \(k^{\frac{1}{k-1}}\) an integer for more than k=2?

OpenStudy (solomonzelman):

Steve, to your initial question \(\color{black}{0\times 0=0+0}\). (\(\color{black}{2}\) is not the only number like this) -------------------------------------- I guess maybe what you noticed is ... \(\tiny \\[0.5em]\) \(\color{black}{2(2^{1/1})=(2^{1/1})^2}\) \(\color{black}{3(3^{1/2})=(3^{1/2})^3}\) \(\color{black}{4(4^{1/3})=(4^{1/3})^4}\) \(\color{black}{5(5^{1/4})=(5^{1/4})^5}\) and so .... \(\color{black}{k\left[k^{\frac{1}{k-1} }\right]=\left[k^{\frac{1}{k-1} }\right]^k}\) for all \(\color{black}{ k \in \mathbb{N}}\).

OpenStudy (solomonzelman):

and actually I like that conclusion, Steve:)

OpenStudy (solomonzelman):

\(\color{red}{\bf Correction:}\) This is true for all integers n\(\ge\)2. (1/0 is undefined)

OpenStudy (solomonzelman):

\(\color{black}{k\left[k^{\frac{1}{k-1} }\right]=\left[k^{\frac{1}{k-1} }\right]^k }\) \(\color{black}{k=\left[k^{\frac{1}{k-1} }\right]^{k-1} }\) \(\color{black}{k=k }\)

OpenStudy (solomonzelman):

this is a brief proof of what you noticed:)

OpenStudy (solomonzelman):

So, we can also assert, by the same token, \(\large \color{black}{k\left[k^{\frac{1}{k-\alpha}}\right]=\left[k^{\frac{1}{k-\alpha}}\right]^{k+1-\alpha } }\)

OpenStudy (solomonzelman):

but it all comes out from k=k.

OpenStudy (mhchen):

Solomon Actually Proved it O_O